# Equivariant Neural Networks for General Linear Symmetries on Lie Algebras

### v1

@@ 5.3.1 Results and Analysis
Figure 4: Qualitative comparison of the best-case (middle), average (left), and challenging (right) test sequences. ReLN models consistently track the ground truth (black) with high fidelity, especially when leveraging covariance. Insets provide a magnified view of the two best-performing variants (ours) to highlight their accuracy.
Our experiments, summarized in Table 5 and Figure 4, reveal a clear performance hierarchy where the geometric representation of features is the critical factor.
Non-equivariant models fail to generalize. Non-equivariant models like ResNets fail to generalize to rotated trajectories, and their performance does not improve when provided with covariance data.
Equivariance is necessary, but implementation design critically affects performance. Both VNs and ReLNs establish strong velocity-only equivariant baselines. ReLNs show improved performance due to its richer feature learning from both the Lie bracket and our bilinear nonlinearity form. However, a method for incorporating covariance is crucial. The VN baseline, which decomposes covariance, degrades the performance. This result confirms that separating the principle axes and its corresponding scales (eigenvalues) prevents the network from learning the structure between a measurement and its uncertainty.
Joint processing of velocity and covariance further improves performance. Integrating covariance as a geometric object in our framework further reduces the ATE, showcasing the performance gains from unified geometric representation on symmetry-preserving uncertainty handling.
Log-covariance ablation highlights mathematical and practical contributions. Finally, the ReLN (Velocity + $\log$-Covariance) model achieves the best performance. By processing covariance matrices via the matrix logarithm, we demonstrate the power of extending Lie-algebraic architectures to operate on manifold-valued data, such as the $\mathrm{SPD}(3)$ matrices representing geometric uncertainty. This success establishes ReLNs as a practical and high-performance framework for uncertainty-aware learning in dynamical systems.
Discussion Our results indicate that ReLNs operate effectively as a geometry- and uncertainty-aware estimator that generalizes across random measurement-frame changes (i.e., arbitrary 3D rotations). While the network is not a classical recursive Markovian filter, it learns to integrate velocity measurements with their associated covariances using uncertainty-dependent weighting, producing robust and accurate trajectory estimates. This behavior suggests ReLNs are suitable as modular components in downstream systems that require handling of matrix-valued uncertainty. Future work will investigate theoretical guarantees for the learned weighting, extensions to other geometric matrix representations, and scalability to higher-dimensional structured geometric inputs.
@@ 6 Conclusion
This work introduces ReLNs, a unified neural architecture that provides exact equivariance to the adjoint action of the general $n\times n$ matrix algebra $\mathfrak{gl}(n)$ and its subgroups. ReLNs enable efficient learning on Lie-algebraic features and structured geometric data, such as covariance matrices. Furthermore, our work establishes a unified Lie-algebraic framework that handles both classical left-action symmetries on vectors and native adjoint-actions on matrices within a single architecture. ReLNs achieve state-of-the-art results on benchmarks and deliver large gains in a challenging drone state estimation task by integrating uncertainty. We’ll apply our equivariant matrix processing capability to a wider array of physical systems, including the dynamics of articulated robots and large-scale sensor fusion, to further expand the boundaries of geometric deep learning.
@@ Appendix A Lie-theoretic preliminaries
This appendix provides an overview of key concepts and derivations from Lie group theory relevant to our construction of $\mathrm{GL}(n)$-equivariant neural networks referenced in the main text.
@@ A.3 Vectorized adjoint
Using the hat/vee maps, the adjoint action on the Lie algebra induces a corresponding action on the vector coordinates. This vectorized action is a linear map represented by a matrix:
@@
In practice we precompute or assemble the $m\times m$ matrix representing $\mathrm{Ad}^{m}_{g}$ (or apply it implicitly) to implement left-multiplicative equivariant layers that act on vector features.
@@ A.4 Structure of Lie Algebra: Semisimplicity and Reductivity
A Lie algebra $\mathfrak{g}$ is:
  • Semisimple if it is a direct sum of simple Lie algebras (i.e., non-abelian and having no nontrivial ideals).
  • Reductive if it decomposes as $\mathfrak{g}=\mathfrak{s}\oplus\mathfrak{z}$, where $\mathfrak{s}$ is semisimple and $\mathfrak{z}$ is the center (an abelian Lie subalgebra).
@@ A Lie algebra $\mathfrak{g}$ is:
$\mathfrak{gl}(n)=\mathfrak{sl}(n)\oplus\mathbb{R}I,$ (14)
where $\mathfrak{sl}(n)$ (traceless matrices) is semisimple, and $\mathbb{R}I$ (scalar matrices) forms the center.
@@
Theorem A.1 (Cartan criterion; standard).
The Killing form $B(X,Y)=\operatorname{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y})$ is non-degenerate iff $\mathfrak{g}$ is semisimple.
In particular, because $\mathfrak{gl}(n)$ contains the central scalar direction $\mathbb{R}I$ with $\mathrm{ad}_{I}=0$, the Killing form is degenerate on $\mathfrak{gl}(n)$.
A.5 Invariant bilinear forms; trace form
Definition A.2.
@@ A.5 Invariant bilinear forms; trace form
$B(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=B(X,Y)\quad\forall g\in G.$ (15)
On semisimple algebras the Killing form provides such an invariant, non-degenerate form. On $\mathfrak{gl}(n)$ we instead use the trace form:
$\langle X,Y\rangle_{\mathrm{tr}}=\operatorname{tr}(XY).$ (16)
Proposition A.1.
The trace form is $\mathrm{Ad}$-invariant on $\mathfrak{gl}(n)$:
$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY).$ (17)
@@
$\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$ by cyclicity of trace. ∎
@@
A.6 Killing form on $\mathfrak{sl}(n)$
Restricted to $\mathfrak{sl}(n)$ the Killing form simplifies and is non-degenerate; one frequently uses the proportionality $B(X,Y)\propto\operatorname{tr}(XY)$ on $\mathfrak{sl}(n)$.
Appendix B Proofs of Key Theorems
B.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$

### v2

@@ 5.2.1 Results and Analysis
Figure 4: Trajectory reconstruction across flight difficulties: high (left), best-case (middle), and medium (right). ReLN models consistently track the ground truth (black) with high fidelity, especially when leveraging covariance. Insets provide a magnified view of the two best-performing variants (ReLN) to highlight their accuracy.
Table 3 summarizes the main findings (full results in Appendix I.6).
Equivariance is necessary for robustness and accuracy. Non-equivariant ResNets exhibit large trajectory errors and do not reliably benefit from adding covariance, indicating that simply concatenating matrix-valued uncertainty does not reliably yield geometry-consistent fusion. Moreover, their errors remain high under test-time $\mathrm{SO}(3)$ rotations, confirming poor robustness to measurement-frame changes.
The Necessity of Structural Compatibility in Covariance Fusion. Among equivariant baselines, velocity-only models already attain low error (e.g., VN), but performance becomes highly sensitive once covariance is introduced. In particular, the VN $(v,C)$ variant collapses toward non-equivariant performance (Table 3; Table 13), consistent with the fact that an eigendecomposition-based interface can separate coupled degrees of freedom (eigenvectors vs. eigenvalues) and yield an unstable or non-geometric uncertainty signal for learning. By contrast, TFN and SE(3)-Transformer can incorporate $(v,C)$ through steerable tensor representations, leading to moderate improvements.
Local equivariant operators offer superior robustness for uncertainty fusion. While SE(3)-Transformer shows a slight performance advantage in the velocity-only regime ($v$), the introduction of covariance uncertainty shifts the advantage to TFN. As shown in Table 13, TFN consistently improves over SE(3)-Transformer across all covariance-integrated variants (e.g., $(v,\log C)$). This pattern suggests that while attention mechanisms can effectively model global dependencies in simple velocity sequences, local convolutional aggregation provides a stronger and more stable inductive bias for processing high-frequency dynamics and uncertainty.
ReLN yields the most consistent uncertainty-aware gains.
ReLN achieves the best overall performance, and its strongest configuration uses $(v,\log C)$ (ATE $13.92$, RTE $11.04$). Unlike tensorial pipelines, ReLN represents velocity and covariance in a shared matrix-algebraic space with a single conjugation rule, enabling uncertainty-conditioned contractions and gating without violating equivariance. The improvement from $C$ to $\log C$ further supports that respecting the intrinsic geometry of $\mathrm{SPD}(3)$ via the matrix logarithm improves conditioning for learning.
Ablative Validation of Reductive Decomposition.
ReLN improves over its semisimple-only variant (the no-center model, equivalent to Lie Neurons), indicating that incorporating the central component $\mathfrak{z}(\mathfrak{g})$ is beneficial for uncertainty-aware estimation. Conversely, removing the semisimple ideal (retaining only the center) degrades performance, increasing ATE from $13.92$ to $16.27$ (Table 3). Together, these ablations support that the best performance arises from modeling the full reductive structure.
Discussion.
Overall, the results indicate that ReLNs act as a geometry- and uncertainty-aware estimator, remaining stable under random measurement-frame changes (test-time $\mathrm{SO}(3)$ rotations). While ReLN is not a classical recursive (Markovian) filter, it learns to fuse velocity sequences with their time-varying covariances through uncertainty-dependent modulation, improving trajectory reconstruction. This suggests ReLNs as a modular backbone for estimation pipelines that must handle matrix-valued uncertainty. A natural direction is to characterize when the learned fusion approximates uncertainty-adaptive integration, and to integrate ReLNs with classical state-estimation frameworks.
5.3 Equivariance for 3D Gaussian Splatting (3DGS)
Figure 5: The ReLN-integrated Gaussian-MAE Framework. Our architecture pipelines raw 3D Gaussian splats through a Geometric Lifting stage to map them into the $\mathfrak{gl}(3)$ Lie algebra. The model explicitly bifurcates the data flow: (1) Active Geometry (position $\mu$, covariance $\Sigma$) is processed via the ReLN Encoder and Decoder to maintain $\mathrm{GL}(3)$-equivariance, while (2) Invariant Attributes (color $c$, opacity $\alpha$) are integrated as Type-0 features. A dedicated Central Positional Embedding ensures stable spatial context. Finally, the Geometric Projector utilizes the Vee map and Killing form to reconstruct the physical Gaussian parameters while preserving holistic symmetry.
Standard point cloud networks typically process only spatial coordinates. However, 3DGS represents scenes using anisotropic 3D Gaussians parameterized by a mean position $\mu\in\mathbb{R}^{3}$ and a covariance matrix $\Sigma\in\mathrm{SPD}(3)$. This introduces a geometric challenge: a global rotation $R\in\mathrm{SO}(3)$ acts differently on these components, where $\mu$ transforms as a vector ($R\mu$) and $\Sigma$ transforms via congruence ($R\Sigma R^{\top}$).
To evaluate ReLNs in this multi-modal geometric setting, we adapt the masked autoencoder framework proposed by Ma et al. (2025), which establishes a self-supervised pretraining baseline for 3D Gaussian Splats. While their original architecture processes Gaussian parameters as loosely coupled features using standard Transformers, we redesign the encoder and decoder using ReLN blocks to enforce $\mathrm{GL}(3)$-equivariance. The overall architecture of our ReLN-integrated Gaussian-MAE is illustrated in Figure 5. Specifically, we unify these geometric types by embedding them into the Lie algebra $\mathfrak{gl}(3)$: the mean position $\mu$ is treated as a translation, while the covariance $\Sigma$ is mapped to the linear subspace $\mathrm{Sym}(3)\subset\mathfrak{gl}(3)$ via the matrix logarithm. This ensures that the network learns consistent geometric representations of 3D shapes under arbitrary orientations.
Experimental Setup. We utilize the ShapeNet dataset processed into 3D Gaussian primitives for self-supervised pre-training, following the protocol of Ma et al. (2025). The model is then fine-tuned on ModelNet10 for classification. We compare our ReLN-based architecture against the baseline model from Ma et al. (2025). Refer to Appendix K for full experimental details.
Experimental Results.
Table 4 evaluates classification accuracy on aligned and randomly rotated ($0^{\circ}$$180^{\circ}$) test sets. The baseline drops from $93.39\%$ to $18.28\%$. In contrast, ReLN maintains a stable accuracy of $95.15\%$, representing a marginal $0.33$ point variance compared to its aligned performance. Furthermore, ReLN converges in fewer epochs during pretraining and achieves lower reconstruction error across attributes (including rotation and scale); see Appendix Figure 8. Together with the rotation-stable accuracy in Table 4, these results support that enforcing the coupled transformation rules of $(\mu,\Sigma)$ via a reductive Lie-algebraic backbone improves robustness to orientation changes.
Table 4: Rotation robustness on 3D Gaussian splats. Classification accuracy (%) on ModelNet10 for the baseline (Ma et al., 2025) and ReLN. Standard uses the aligned test set; Rotated applies random rotations ($0^{\circ}$$180^{\circ}$).
Method Standard Acc. (%) Rotated Acc. (%)
Baseline (Ma et al., 2025) 93.39 18.28
ReLN (Ours) 94.82 95.15
5.4 EMLP Double-Pendulum Benchmark and Efficiency
We evaluate ReLN on the EMLP double-pendulum benchmark Finzi et al. (2021), which learns Hamiltonian dynamics under multiple symmetry groups ($O(2)$, $SO(2)$, and $D_{6}$). Following the standard protocol, we report test rollout errors in Table 5; additional details are provided in Appendix L.
Accuracy. ReLN achieves comparable or lower rollout error than EMLP across all symmetry settings (e.g., $SO(2)$: 0.010 vs. $0.015$), without introducing group-specific architectural modifications.
Compute and latency. With matched hidden size and representation dimension, ReLN-HNN is substantially cheaper per step than EMLP-HNN (Table 6): 142,190142{,}190 vs. 1,589,9091{,}589{,}909 FLOPs ($11.18\times$) and $2.216$ vs. $61.635$ ms ($27.81\times$). This gap is consistent with ReLN using closed-form matrix operations from exact $\mathrm{Ad}$-equivariant primitives on $\mathfrak{gl}(n)$, whereas EMLP relies on symmetry-dependent basis construction and projection.
Overall, these results support ReLN as a reusable equivariant backbone that attains EMLP-level accuracy while substantially reducing per-step computation across different symmetry choices.
Table 5: Test rollout error on the EMLP Double-Pendulum Benchmark. ReLN matches or improves upon EMLP across different symmetry groups ($O(2)$, $SO(2)$, $D_{6}$). Values represent the geometric mean of relative errors, with the standard deviation over 3 trials in parentheses; e.g., the notation $0.012(2)$ denotes $0.012\pm 0.002$.
Metric $O(2)$ $SO(2)$ $D_{6}$ MLP-HNNs
EMLP ReLN EMLP ReLN EMLP ReLN
Rollout Error 0.012 (2) 0.011 (2) 0.015 (3) 0.010 (4) 0.013 (2) 0.011 (2) 0.028
Table 6: Computational cost on the HNN task. ReLN-HNN reduces FLOPs/step and inference latency relative to EMLP-HNN.
Model # Params FLOPs / step Inference (ms)
MLP-HNN 34,817 70,400 0.159
EMLP-HNN 55,569 1,589,909 61.635
ReLN-HNN (Ours) 69,889 142,190 2.216
6 Discussion and Limitations.
ReLNs are exactly $\mathrm{GL}(n)$-adjoint equivariant by construction. For covariance-like quantities, however, the physically standard transformation is congruence, $\Sigma\mapsto A\,\Sigma\,A^{\top}$, which agrees with adjoint conjugation $X\mapsto AXA^{-1}$ only for orthogonal changes of frame ($A^{-1}=A^{\top}$). Thus, while our unified Lie-algebraic interface is most directly interpretable under rigid rotations, extending the same measurement–uncertainty coupling to general $\mathrm{GL}(n)$ coordinate changes requires additional geometric justification and may call for alternative liftings or invariants. A further direction is to extend the framework to groups with translations, such as affine and $\mathrm{SE}(3)$ symmetries, by leveraging their semidirect-product structure and the induced actions on the corresponding Lie algebras.
@@ 7 Conclusion
We introduced Reductive Lie Neurons (ReLNs), a unified architecture that is exactly equivariant to the adjoint action on $\mathfrak{gl}(n)$ and its reductive matrix subalgebras. Using a non-degenerate $\mathrm{Ad}$-invariant contraction, ReLNs provide an efficient Lie-equivariant framework for learning with structured geometric quantities, including matrix-valued uncertainty. The same viewpoint links classical left actions on vectors to adjoint actions on matrices via explicit Lie-algebraic liftings, enabling a single backbone to process heterogeneous geometric inputs in a consistent manner. Empirically, ReLNs achieve strong performance on algebraic benchmarks and 3D Gaussian-splat representation learning, and yield substantial gains on drone state estimation by incorporating uncertainty equivariantly.
We expect these capabilities to benefit data-efficient representation learning in vision and robotics, particularly for structured 3D scene models—such as 3D Gaussian splats and world models—where coupled transformation rules are central. Future work will scale ReLNs to larger physical systems (e.g., multimodal fusion and articulated dynamics) and further explore efficient equivariant architectures for matrix-parameterized representations.
@@ Appendix A Lie-theoretic preliminaries
This appendix provides an overview of key concepts and derivations from Lie group theory relevant to our construction of $\mathrm{GL}(n)$ adjoint-equivariant neural networks.
@@ A.3 Vectorized adjoint
Using the hat ($\wedge$) / vee ($\vee$) maps, the adjoint action on the Lie algebra induces a corresponding action on the vector coordinates. This vectorized action is a linear map represented by a matrix:
@@
Equivalently, $\mathrm{Ad}^{m}_{g}$ is the $m\times m$ matrix representation of $\mathrm{Ad}_{g}$ in the chosen basis $\{E_{i}\}$. In practice we precompute or assemble the $m\times m$ matrix representing $\mathrm{Ad}^{m}_{g}$ (or apply it implicitly) to implement left-multiplicative equivariant layers that act on vector features.
@@ A.4 Structure of Lie Algebra: Semisimplicity and Reductivity
A finite-dimensional Lie algebra $\mathfrak{g}$ is
  • semisimple if it has no nonzero solvable ideals (equivalently, its radical is zero);
  • reductive if it decomposes as a direct sum of ideals
    $\mathfrak{g}\;=\;\mathfrak{z}(\mathfrak{g})\;\oplus\;[\mathfrak{g},\mathfrak{g}],$ (14)
    where $\mathfrak{z}(\mathfrak{g})$ is the center and $\mathfrak{g},\mathfrak{g}]$ is semisimple.
@@
$\mathfrak{gl}(n)=\mathbb{R}I\oplus\mathfrak{sl}(n),\qquad\mathbb{R}I=\mathfrak{z}(\mathfrak{gl}(n)),\quad\mathfrak{sl}(n)=[\mathfrak{gl}(n),\mathfrak{gl}(n)].$ (15)
where $\mathfrak{sl}(n)$ (traceless matrices) is semisimple, and $\mathbb{R}I$ (scalar matrices) forms the center. Thus $\mathfrak{gl}(n)$ is reductive but not semisimple (it has a nontrivial center).
@@
The Killing form on $\mathfrak{g}$ is
$B_{\mathfrak{g}}(X,Y):=\operatorname{tr}(\mathrm{ad}_{X}\circ\mathrm{ad}_{Y}).$ (16)
Theorem A.2 (Cartan criterion).
For a finite-dimensional Lie algebra $\mathfrak{g}$, the Killing form $B_{\mathfrak{g}}$ is non-degenerate if and only if $\mathfrak{g}$ is semisimple.
In particular, if $\mathfrak{g}$ has nontrivial center then $B_{\mathfrak{g}}$ is degenerate, since $\mathrm{ad}_{Z}=0$ for all $Z\in\mathfrak{z}(\mathfrak{g})$.
A.5 $\mathrm{Ad}$-invariant bilinear forms and the trace form
Definition A.3 ($\mathrm{Ad}$-invariance).
@@ A.5 $\mathrm{Ad}$-invariant bilinear forms and the trace form
$B(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=B(X,Y)\qquad\forall g\in G,\;\forall X,Y\in\mathfrak{g}.$ (17)
On semisimple $\mathfrak{g}$, the Killing form is $\mathrm{Ad}$-invariant and non-degenerate. On reductive but non-semisimple algebras (notably $\mathfrak{gl}(n)$), one often uses alternative $\mathrm{Ad}$-invariant forms. A basic example on $\mathfrak{gl}(n)$ is the trace pairing
$\langle X,Y\rangle_{\mathrm{tr}}:=\operatorname{tr}(XY).$ (18)
Proposition A.4 ($\mathrm{Ad}$-invariance of the trace pairing on $\mathfrak{gl}(n)$).
For $g\in\mathrm{GL}(n)$ and $X,Y\in\mathfrak{gl}(n)$,
$\operatorname{tr}\bigl((gXg^{-1})(gYg^{-1})\bigr)=\operatorname{tr}(XY).$ (19)
@@
By cyclicity of trace, $\operatorname{tr}((gXg^{-1})(gYg^{-1}))=\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$. ∎
@@
We will use explicit $\mathfrak{gl}(n)$ identities (including the closed form of $B_{\mathfrak{gl}(n)}$, the resulting degeneracy, and our non-degenerate $\mathrm{Ad}$-invariant modification $\widetilde{B}$) in Appendix B.
Appendix B Connections for $\mathfrak{gl}(n)$ and $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$
This appendix collects two specialization facts used in the main text: (i) an explicit expression for the (degenerate) Killing form on $\mathfrak{gl}(n)$ and its relation to our modified form $\widetilde{B}$; (ii) the standard intertwining between the $\mathrm{SO}(3)$ vector action and the adjoint action on $\mathfrak{so}(3)$.
B.1 The classical Killing form on $\mathfrak{gl}(n)$
Proposition B.1 (Closed form and degeneracy).
For $X,Y\in\mathfrak{gl}(n)$,
$B_{\mathfrak{gl}(n)}(X,Y)=2n\,\operatorname{tr}(XY)-2\,\operatorname{tr}(X)\operatorname{tr}(Y).$ (20)
In particular, $B_{\mathfrak{gl}(n)}$ is degenerate, and $\mathbb{R}I$ lies in its radical.
Proof sketch.
The identity equation 20 is standard for $\mathfrak{gl}(n)$ (up to an overall normalization convention). Degeneracy follows since $\mathrm{ad}_{\lambda I}=0$ for all $\lambda I\in\mathbb{R}I$, hence $B_{\mathfrak{gl}(n)}(\lambda I,\cdot)=\operatorname{tr}(\mathrm{ad}_{\lambda I}\circ\mathrm{ad}_{\cdot})=0$. ∎
Proposition B.2 (Restriction to $\mathfrak{sl}(n)$).
If $X,Y\in\mathfrak{sl}(n)$, then
$B_{\mathfrak{gl}(n)}(X,Y)=2n\,\operatorname{tr}(XY),$ (21)
which coincides with the Killing form on the semisimple ideal $\mathfrak{sl}(n)$ under the same normalization.
Proof.
For $X,Y\in\mathfrak{sl}(n)$, $\operatorname{tr}(X)=\operatorname{tr}(Y)=0$ and equation 20 reduces to $2n\,\operatorname{tr}(XY)$. ∎
B.2 Our modified $\mathrm{Ad}$-invariant form on $\mathfrak{gl}(n)$
In the main text we use
$\widetilde{B}(X,Y)=2n\,\operatorname{tr}(XY)-\operatorname{tr}(X)\operatorname{tr}(Y).$ (22)
Proposition B.3 (Relation to $B_{\mathfrak{gl}(n)}$).
With $B_{\mathfrak{gl}(n)}$ as in equation 20,
$\widetilde{B}(X,Y)=B_{\mathfrak{gl}(n)}(X,Y)+\operatorname{tr}(X)\operatorname{tr}(Y).$ (23)
Proof.
Subtract equation 20 from equation 22. ∎
Proposition B.4 (Orthogonal splitting).
Write $X=X_{0}+\frac{1}{n}\operatorname{tr}(X)I$ and $Y=Y_{0}+\frac{1}{n}\operatorname{tr}(Y)I$ with $X_{0},Y_{0}\in\mathfrak{sl}(n)$. Then
$\widetilde{B}(X,Y)=2n\,\operatorname{tr}(X_{0}Y_{0})+\operatorname{tr}(X)\operatorname{tr}(Y),$ (24)
and $\mathbb{R}I$ and $\mathfrak{sl}(n)$ are $\widetilde{B}-$orthogonal.
Proof.
Using $\operatorname{tr}(I)=n$ and $\operatorname{tr}(X_{0})=\operatorname{tr}(Y_{0})=0$,
$\operatorname{tr}(XY)=\operatorname{tr}(X_{0}Y_{0})+\frac{1}{n}\operatorname{tr}(X)\operatorname{tr}(Y).$ (25)
Substitute into equation 22. Orthogonality follows by setting $X=\lambda I$ and $Y\in\mathfrak{sl}(n)$. ∎
Proposition B.5 ($\mathrm{Ad}$-invariance).
For all $g\in\mathrm{GL}(n)$ and $X,Y\in\mathfrak{gl}(n)$,
$\widetilde{B}(\mathrm{Ad}_{g}X,\mathrm{Ad}_{g}Y)=\widetilde{B}(X,Y).$ (26)
Proof.
Use $\mathrm{Ad}_{g}X=gXg^{-1}$, cyclicity $\operatorname{tr}(gXYg^{-1})=\operatorname{tr}(XY)$, and $\operatorname{tr}(gXg^{-1})=\operatorname{tr}(X)$ in equation 22. ∎
Proposition B.6 ($\widetilde{B}$ as an $\mathrm{Ad}$-invariant non-degenerate completion on the center).
The classical Killing form $B_{\mathfrak{gl}(n)}$ is degenerate with radical containing the center $\mathfrak{z}(\mathfrak{gl}(n))=\mathbb{R}I$. The modified form $\widetilde{B}$ in equation 22 is $\mathrm{Ad}$-invariant and non-degenerate, and it agrees with $B_{\mathfrak{gl}(n)}$ on the semisimple ideal $\mathfrak{sl}(n)$ while supplying a non-degenerate inner product on the center:
$\widetilde{B}|_{\mathfrak{sl}(n)\times\mathfrak{sl}(n)}=B_{\mathfrak{sl}(n)},\qquad\widetilde{B}(\lambda I,\mu I)=n^{2}\,\lambda\mu\quad(\lambda,\mu\in\mathbb{R}).$ (27)
In particular, $\widetilde{B}$ resolves the degeneracy of $B_{\mathfrak{gl}(n)}$ precisely along $\mathbb{R}I$ without altering the semisimple part.
Proof.
$\mathrm{Ad}$-invariance is Proposition B.5. For $\lambda I,\mu I\in\mathbb{R}I$, equation 22 gives $\widetilde{B}(\lambda I,\mu I)=2n\,\operatorname{Tr}(\lambda\mu I)-\operatorname{Tr}(\lambda I)\operatorname{Tr}(\mu I)=2n(\lambda\mu n)-(\lambda n)(\mu n)=n^{2}\lambda\mu$, which is non-degenerate. The restriction to $\mathfrak{sl}(n)$ follows since $\operatorname{Tr}(X)=0$ on $\mathfrak{sl}(n)$, so equation 22 reduces to $2n\,\operatorname{Tr}(XY)$, which matches Proposition B.2 under the same normalization. ∎
B.3 $\mathfrak{so}(3)\simeq\mathbb{R}^{3}$: vector action vs. adjoint action
Let $\hat{\cdot}:\mathbb{R}^{3}\to\mathfrak{so}(3)$ be the standard hat map defined by $\hat{v}\,w=v\times w$, with inverse $(\cdot)^{\vee}:\mathfrak{so}(3)\to\mathbb{R}^{3}$.
Lemma B.7 (Intertwining identity).
For all $R\in\mathrm{SO}(3)$ and $v\in\mathbb{R}^{3}$,
$\widehat{Rv}=R\hat{v}R^{\top}=\mathrm{Ad}_{R}(\hat{v}),\qquad\mathrm{Ad}_{R}(X)=RXR^{-1}=RXR^{\top}.$ (28)
[…]
Proof.
For any $w\in\mathbb{R}^{3}$, using $(Ra)\times(Rb)=R(a\times b)$,
$(R\hat{v}R^{\top})w=R\hat{v}(R^{\top}w)=R\bigl(v\times(R^{\top}w)\bigr)=(Rv)\times w=\widehat{Rv}\,w.$ (29)
Thus $\widehat{Rv}=R\hat{v}R^{\top}$. The equivariance statement follows by applying $(\cdot)^{\vee}$ and using $\widehat{Ra}=R\hat{a}R^{\top}$. ∎
Remark B.8 (Trace contraction recovers the Vector Neuron inner product).
Under the hat identification $\hat{\cdot}:\mathbb{R}^{3}\to\mathfrak{so}(3)$,
$\hat{v}=\begin{bmatrix}0&-v_{3}&v_{2}\\ v_{3}&0&-v_{1}\\ -v_{2}&v_{1}&0\end{bmatrix},\qquad\operatorname{Tr}(\hat{v}\,\hat{w})=-2\,v^{\top}w\quad(v,w\in\mathbb{R}^{3}).$ (30)
Hence the $\mathrm{Ad}$-invariant trace pairing on $\mathfrak{so}(3)$, $\langle X,Y\rangle_{\mathrm{tr}}:=-\tfrac{1}{2}\operatorname{Tr}(XY)$, corresponds exactly to the Euclidean dot product on $\mathbb{R}^{3}$: $\langle\hat{v},\hat{w}\rangle_{\mathrm{tr}}=v^{\top}w$. In particular, $\mathrm{Ad}$-invariant scalar contractions on $\mathfrak{so}(3)$ recover the invariants used by Vector Neurons  Deng et al. (2021)(up to a fixed global scaling, which can be absorbed by adjacent learnable linear maps or normalization).
Appendix C Proofs of Key Theorems
In this section, we provide a generalized treatment and proofs for the modified bilinear form $\widetilde{B}$ introduced in Section 4 and Appendix B. Our discussion here establishes the formal properties of $\widetilde{B}$ for any real reductive Lie algebra.
C.1 Proof of Non-degeneracy and Ad-invariance of Modified Killing Form $B_{e}$